• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Deur A.

    Log entry time 16:03:33 on October 25,1998

    Entry number 3589

    keyword=Shift summary

    E=3.4 GeV, P=3.028 GeV/c, target in transverse configuration. Production runs
    8:50 Asked MCC for Bscope+checked raster size. Good
    8:56 target Ion chamber readout: 106.2 Rad (total dose)
    12:00 NMR measurement (have to switch from |_ to \\ , then NMR, then back to |_. Target polarization 28 % (loss). Used new calibration from Gordon
    14:07 coda crashed (100 % deadtime, 0 Hz)]. It occured after we stop the last run because the current disc was full.
    16:00 beam down (tripped after have secure the beam for hall B)
    Total number of events so far. 48.5 M over 85 M